Database last updated on Sun, 01 Feb 2026 14:40:50 +0000 by localhost

Critical warnings from b-files' comments for days between 2024066 and 2024066 (0 sec, 6 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
rl:2024 066 16 24 28 Motor #3: appears stuck with a reference sensor activated.21#19405:30:531-1
rl:2024 066 16 25 32 New value set for watermark SUPPLY.VOLTAGE.HIGH[133#19405:30:581-1
rl:2024 066 01 33 40 PMT counter failure11#19405:31:421-1
uvfocusd:Lamp not on... uvfocusd test terminated F1= 2 F4= 251#19416:46:191-1
hg:Lamp not detected f1= 2 f4= 211#19420:13:001-1
skc:Azimuth drive turned OFF in op_st.12#19421:28:331-1

Instrument constant changes for the week before 2024066 (5 sec, 0 rows)

HG burned out? (AP results when HG is ON) (0 sec, 13 rows)

SerialYYYYJJJHG CurrentHG Voltage
19420240590.1813.83
19420240600.1813.84
19420240610.1813.83
19420240620.1813.83
19420240630.1813.8
19420240640.1813.81
19420240650.1813.8
19420240650.1813.81
19420240650.1813.78
19420240660.1813.81
19420240660.1813.78
19420240660.1813.8
19420240660.1813.75

Hg with step corrections larger than 10 steps or multiple corrections larger than 2.5 steps (0 sec, 1 rows)

SerialYYYYJJJHG largemax correction
19420240661818.3

HP with step corrections larger than 20 steps (0 sec, 1 rows)

SerialYYYYJJJHP largemax correction
1942024066432.6

Non-critical warnings from b-files' comments for days between 2024066 and 2024066 (0 sec, 8 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
rl:2024 066 16 24 27 Checksum bad in low level command.13#19405:30:501-1
rl:2024 066 16 24 28 Motor #3: appears stuck with a reference sensor activated.13#19405:30:531-1
rl:2024 066 16 24 28 Motor #11: appears stuck with a reference sensor activated.13#19405:30:551-1
pd:Going to menu299#19412:36:031-1
dsp:Doing dispersion test with Cd (UV port)299#19413:24:311-1
checkhg:Last HG had to adjust the micrometer by at least 1299#19414:42:491-1
uvfcousr:Illegal command (uvfcousr), number 4/ 4, in schedule , ignoring.199#19418:18:101-1
dp:Doing dispersion with internal HG.199#19422:31:581-1

Measurements and tests summary.


_jday_loc_serial_skcaodedsdtedhghprssl
066Toronto#194Azimuth drive turned OFF in op_st.,calsc,menu,Saving backup OP_ST.000,Starting schedule inside,Waiting until ZA = 170 now 128.1,Waiting until ZA = 170 now 129.5,Waiting until ZA = 170 now 130.9,Waiting until ZA = 170 now 132.2,Waiting until ZA = 170 now 133.5,Waiting until ZA = 170 now 134.6,Waiting until ZA = 170 now 135.8,Waiting until ZA = 170 now 136.8,Waiting until ZA = 170 now 137.7,Waiting until ZA = 170 now 138.6,Waiting until ZA = 170 now 139.3,Waiting until ZA = 170 now 139.9,Waiting until ZA = 170 now 140.4,Waiting until ZA = 170 now 140.8,Waiting until ZA = 170 now 141.1,Waiting until ZA = 170 now 141.24452535255
(0 sec)

Relative and absolute humidity for day 2024066 (1 sec, 1 rows)

SerialRH, %AH, g/m37-day RH historyDes-nt replcd days ago
19428.06.8